# An End-to-End Depth-Based Pipeline for Selfie Image Rectification

### v1

@@ Abstract
Portraits or selfie images taken from a close distance typically suffer from perspective distortion. In this paper, we propose an end-to-end deep learning-based rectification pipeline to mitigate the effects of perspective distortion. We learn to predict the facial depth by training a deep CNN. The estimated depth is utilized to adjust the camera-to-subject distance by moving the camera farther, increasing the camera focal length, and reprojecting the 3D image features to the new perspective. The reprojected features are then fed to an inpainting module to fill in the missing pixels. We leverage a differentiable renderer to enable end-to-end training of our depth estimation and feature extraction nets to improve the rectified outputs. To boost the results of the inpainting module, we incorporate an auxiliary module to predict the horizontal movement of the camera which decreases the area that requires hallucination of challenging face parts such as ears. Unlike previous works, we process the full-frame input image at once without cropping the subject’s face and processing it separately from the rest of the body, eliminating the need for complex post-processing steps to attach the face back to the subject’s body.
To train our network, we utilize the popular game engine Unreal Engine to generate a large synthetic face dataset containing various subjects, head poses, expressions, eyewear, clothes, and lighting.
Quantitative and qualitative results show that our rectification pipeline outperforms previous methods, and produces comparable results with a time-consuming 3D GAN-based method while being more than 260 times faster.
@@ 1 Introduction
  • We will release the source code, trained models, and the dataset upon the publication of the paper.
2 Related Work
@@
The prior works that directly address the perspective distortion of close-range portrait can be categorized into 3 classes: Conventional optimization-based [6], learning-based [1], and 3D GAN-based methods [7]. Fried [6] adjusts the distance between the camera and subject given a single portrait by estimating an initial head model and camera intrinsic and extrinsic parameters where they are jointly optimized to find the optimal head model, expression and camera parameters that fit pre-detected facial landmarks. This 3D model is then used to compute a 2D smooth, pixel-wise motion field which produces no holes after warping the original image.
@@
Regarding the training dataset, LPUP [8] synthesizes training samples from a static 3D head model captured by a light-stage system. However, this model has fixed expressions, head orientation and limited representation of the subject’s hair. To increase the number of samples, they utilize a publicly available dataset of faces with their corresponding depth maps [9]. This dataset also has limited head orientation and mainly captures the subject looking forward to the camera along with fixed expression and illumination. The distorted images were then rendered simply from the RGBD images using perspective projection and rasterization, without considering changes in illumination or complex reflectance properties [8]. In contrast, we leverage the popular game engine Unreal Engine (UE), to generate photorealistic image pairs with various expressions, eyeglasses, grooms, head orientation, camera-head poses, illumination and body clothes.
Recently, Disco [7] was proposed to adjust the camera-to-subject distance with the help of 3D GANs. They rely on the recent powerful generation capability of a pretrained 3D GAN, known as EG3D [14], to generate plausible face corrections. The 3D GANs are a type of generative network that is additionally conditioned on camera parameters to generate appearance-consistent facial images under different view angles [14, 15]. To manipulate the camera-to-subject distance, they first estimate an initial head model and camera pose and intrinsics that are jointly optimized through the GAN inversion technique [16] using a combination of the photometric-based loss LPIPS [17] and the L2 loss between the landmark detected on the generated and input face to find the camera parameters, face identity and expressions (i.e. the latent face code) that when provided to the generator produces the distorted input image. The generator is then finetuned (overfit) on this selfie input until convergence. The inversion process is time-consuming as the authors of Disco [7] report that this operation requires more than 2 minutes on a modern GPU. In addition, their pipeline produces subtle-to-noticeable identity changes, depending on the input, due to the inaccuracies of the encoded face from the inversion of 3D GAN outputs. Moreover, similar to [8], their methods require cropping the face from the body leading to additional complex post-processing steps to combine the face with the rest of the body and the background image afterwards.
@@
The work of [19], was not mainly addressing the task of rectifying perspective distortion, but they were performing it as a subtask of their pipeline which aims to synthesize a selfie photo of the whole body from a group of close-range shots of different body parts. As the face forms a small portion of the final synthesized image, the quality of the corrected face was not the main target of the method. They follow a flow-based warping method similar to LPUP [8] and train the architecture from [20] on image pairs generated by the 3D GAN network EG3D [14]. However, the artefacts in the generated training data led to inferior results compared to previous methods.
@@ 3.1 Overview
3.2 Depth Prediction
We use a residual UNet architecture similar to many recent monocular depth estimation methods [8] for facial depth prediction. The encoder and decoder consist of 5 residual blocks for each. The input to the depth network is the segmented foreground human. Hence the network only estimates the depth over the foreground region.
3.3 Feature Extraction
Instead of directly reprojecting the image in the RGB space, we extract high-level features from the foreground-segmented input image with another UNet-based architecture composed of 5 mirrored pairs of convolution layers. The decoder outputs a feature map that has the same spatial size as the input but with 64 channels. Thus, every depth pixel is attached to a richer feature vector of length 64.
3.4 Image Warping
Once we have the depth map of the subject, we are able to synthesize a novel view of the subject by manipulating the camera-to-subject distance and the focal length. We unproject the face to the 3D space of the original camera $C_{1}$ that has camera intrinsics matrix $K_{1}$ expressed as the following:
@@ 3.4 Image Warping
Rendering the actual image (i.e. warping the image) is called the rasterization process. To enable E2E training of our model, we leverage a differentiable renderer module of PyTorch3D framework [5].
@@
3.5 Camera Horizontal Translation
@@
To compute the rotation $\theta$, we choose an arbitrary point in the face where this point after translating and rotating the camera by $t_{x}$ and $\theta$ respectively, should have the same position as if we moved the camera back with $t_{x}$ and $\theta$ are zeros. We choose this pixel point on the left face boundary if the subject is closer to the left image side and vice versa. Choosing the point on face boundaries guarantees that this point will not go outside the image boundary after the reprojection. This is especially important when the face is close to the image boundaries; otherwise, any point is valid. Let’s suppose the chosen pixel point $p$ in the input image is at position $(u,v)$ expressed in 2D image coordinates. The 3D coordinates of the point $p$ in camera $C_{2}$ is $P_{2}$ = $[X_{2},Y_{2},Z_{2}]^{T}$. The transformation of $P_{2}$ to the coordinates of camera $C^{\prime}_{2}$ is:
@@
We use a ResNet18 [23] architecture for our translation module. It predicts a vector $V$ of length $l=50$ which corresponds to the equal-spacing discretization of the translation range $[-20,20]$ cm. During inference, we find the translation by computing the Softmax of the output vector as the following:
@@ 3.6 Image Refinement
3.7 Post Processing and Image Composition
@@
3.8 Training and Loss Functions
Depth Loss. To train the depth network, we use a loss function that is typically used in training monocular depth estimation networks [8]. Specifically, we use L1 loss (i.e. the mean absolute error) $\mathcal{L}_{L1}$ between the predicted depth and ground truth depth. We also compute $\mathcal{L}_{tanh}$ which is the L1 loss between the hyperbolic tangent (i.e. $tanh$ function) of each of them, and $\mathcal{L}_{grad}$ which is the multi-scale gradients loss [8]. The overall loss for depth training is then defined as the following:
@@ 3.8 Training and Loss Functions
Horizontal Translation Module Training and Loss. During our experiments, we found it hard to train the translation network (HzT net) in an unsupervised way by incorporating it directly into the pipeline. Thus, we generated the training GT targets to train the network in a supervised way. To compute the GT translations, we discretize the possible range of camera horizontal translation, as mentioned in Section 3.5, to 50 translation steps or bins. For each step, we compute the corresponding rotation according to (6), warp the image to the new pose and provide it to our pretrained generator module which predicts the rectified image and mask as explained in Section 3.6. We compare the predicted output mask $m_{g}$ with the GT target mask $m_{t}$. The details of synthesizing $m_{t}$ are explained in Section 4. We use the intersection over union (IoU) to compute the difference between $m_{g}$ and $m_{t}$ as the following:
@@
We use the binary cross entropy loss between the predicted vector $\hat{V}$ from the HzT net and the target vector $V$. The predicted translation is estimated according to (7) and used to warp the input foreground mask $m$. We then compute the IoU loss between the warped mask $m_{w}$ and $m_{t}$. The total loss for training the HzT module becomes as the following:
@@
4 Dataset Generation
To train our model, we need pairs of perspective-distorted images and their distortion-free targets. In [1], they used BU-4DFE facial dataset [9] for generating training pairs. This dataset has static illumination and expressions for each of the subjects. The 3D head is reconstructed from the synchronized multiple cameras. The reconstruction does not include the subject body and usually has limited variations of head poses. The distorted version of the faces are rendered simply using perspective projection and rasterization, without changing illumination or considering complex reflectance properties. Unlike [1], we utilize Unreal Engine (UE) for rendering a realistic synthetic facial dataset. UE offers realistic human characters called Metahuman (MH). Those MH are captured in different locations of a simulated city project in UE called the City Sample. During capturing, we also vary the camera-to-subject distance, the camera orientations, the head pitch, yaw, roll and facial expressions.
@@ 4 Dataset Generation
We used two types of MH characters: the first is a predefined set of characters provided by UE, and the second is a subclass instance of the MH character called ‘CitysSampleCrowd’ character. The latter is a pool of different faces, outfits, skin texture, and grooms (i.e. hair, beard, moustache, etc). The character is then randomly generated from those various options. In the first type, we used 66 different characters, from which 10 characters are dedicated for testing. To simulate individuals wearing eyeglasses, we utilize a UE asset package containing 18 different types of sunglasses and reading glasses for both men and women.
We generated two sets of synthetic data: a single-view and a multi-view dataset. In the former, the close range and its corresponding far-range images are generated. In the multi-view case, we capture multiple far-range views. Specifically, we generate 4 rear views: left, right, centre, and top side of the front camera, as illustrated in Figure 4. The poses of those cameras are fixed relative to the front camera. This dataset is used for the training and evaluation of the horizontal translation module as explained in Section 3.8.
@@
To train the Horizontal translation module described in Section 3.5, we need to synthesize the ground-truth target image on the fly during training since we have only the view that corresponds to $t_{x}=0$ (i.e. the camera moves straight to the back). To synthesize the target image with $t_{x}\neq 0$, we reproject the view of the rear centre camera to the target pose. Similarly, we warp the other rear views to the same target pose. The missing pixels from the warped centre view are filled from the other reprojected rear views. The synthesized target mask is used during training, while the target colour photo is only used to report the photometric errors.
5 Experiments and Results
5.1 Training details
@@
Next, we pretrain the feature extractor ($E$) and generator ($G$) nets shown in Figure 1 where the ground truth depth is used in projection instead of the predicted one. The loss function is computed according to (11) where $\mathcal{L}_{depth}$ is disabled, and $\omega_{all}$, $\omega_{face}$, $\omega_{missing}$, $\beta$ and $\rho$ are set to $0.2$, $1$, $5$, $1$ and $20$, respectively. Adam optimizer [28] is used and the learning rate is $5\times 10^{-4}$ and halved every 50 epochs with a total epochs of 200.
@@ 5.2 Evaluation metrics
  • Euclidean distance between the estimated facial landmarks of the predicted output and the reference image. We use Mediapipe [30] to estimate dense facial landmarks and use them to align both images by a similarity transform. We then compute the Euclidian distance between the corresponding landmarks.
@@
5.3 Results
We compare our pipeline to the following previous portrait correction solutions: Fried et al. [6], LPUP [1] and Disco [7]. Additionally, we compare our method to Shih’s method [4] which is a solution for correcting a different type of face distortion that occurs in images captured with wide-angle cameras when the subject is close to the corners of the image, which is solved by stereographic projection. Since LPUP’s authors [1] do not release their code or datasets, we reimplemented the first part of the LPUP pipeline so that we can compare their proposed pipeline with ours. Specifically, we replicate the training of the Flownet part of the pipeline. The input to the network consists of an RGB image and a depth label. The depth label is repeated to have the same spatial size as the input image and concatenated with the image channels. The network output is a two-channel deformation map used to warp the input image to correct the perspective distortion. Second, we add a differentiable bilinear sampling module to make a backward warping of the GT distortion-free image using the predicted and ground-truth flow maps to the perspective of the input selfie image. The warped images are labelled as fake and real, respectively, and used to train a discriminator network. Thus, the loss function for training the flow map network becomes similar to LPUP [1]. For a fair comparison, we use our synthetically generated dataset for the training. The dataset is preprocessed in the same way as in LPUP. Particularly, we scale and crop the image such that all faces in the images have approximately the same scale, and match the inner corner of each subject’s right eye to a fixed position through all images.
In the qualitative and quantitative comparison with Fried [6] and Disco [7], we are constrained by the results reported in their paper or project websites, as no codes have been released by both methods at the time of paper submission.

### v2

@@ Abstract
Portraits or selfie images taken from a close distance typically suffer from perspective distortion. In this paper, we propose an end-to-end deep learning-based rectification pipeline to mitigate the effects of perspective distortion. We learn to predict the facial depth by training a deep CNN. The estimated depth is utilized to adjust the camera-to-subject distance by moving the camera farther, increasing the camera focal length, and reprojecting the 3D image features to the new perspective. The reprojected features are then fed to an inpainting module to fill in the missing pixels. We leverage a differentiable renderer to enable end-to-end training of our depth estimation and feature extraction nets to improve the rectified outputs. To boost the results of the inpainting module, we incorporate an auxiliary module to predict the horizontal movement of the camera which decreases the area that requires hallucination of challenging face parts such as ears. Unlike previous works, we process the full-frame input image at once without cropping the subject’s face and processing it separately from the rest of the body, eliminating the need for complex post-processing steps to attach the face back to the subject’s body. To train our network, we utilize the popular game engine Unreal Engine to generate a large synthetic face dataset containing various subjects, head poses, expressions, eyewear, clothes, and lighting. Quantitative and qualitative results show that our rectification pipeline outperforms previous methods, and produces comparable results with a time-consuming 3D GAN-based method while being more than 260 times faster.
@@ I Introduction
II Related Work
@@
The prior works that directly address the perspective distortion of close-range portrait can be categorized into 3 classes: Conventional optimization-based [6], learning-based [1, 14], and 3D GAN-based methods [7]. Fried [6] adjusts the distance between the camera and subject given a single portrait by estimating an initial head model and camera intrinsic and extrinsic parameters where they are jointly optimized to find the optimal head model, expression and camera parameters that fit pre-detected facial landmarks. This 3D model is then used to compute a 2D smooth, pixel-wise motion field which produces no holes after warping the original image.
@@
Regarding the training dataset, LPUP [8] synthesizes training samples from a static 3D head model captured by a light-stage system. However, this model has fixed expressions, head orientation and limited representation of the subject’s hair. To increase the number of samples, they utilize a publicly available dataset of faces with their corresponding depth maps [9]. This dataset also has limited head orientation and mainly captures the subject looking forward to the camera along with fixed expression and illumination. The distorted images were then rendered simply from the RGBD images using perspective projection and rasterization, without considering changes in illumination or complex reflectance properties [8]. In contrast, we leverage the popular game engine Unreal Engine (UE) to generate photorealistic image pairs with various expressions, eyeglasses, grooms, head orientation, camera-head poses, illumination and body clothes.
Deep Face Normalization (DFN) [14] proposed a module to rectify the perspective distortion as part of a pipeline aimed at ’normalizing’ facial appearances. While their module shares similarities with LPUP [1] in that it also generates 2D motion maps, DFN adopts a simpler approach. Specifically, it consists of a single network that takes the distorted face image along with its 2D facial landmarks as input.
Recently, Disco [7] was proposed to adjust the camera-to-subject distance with the help of 3D GANs. They rely on the recent powerful generation capability of a pretrained 3D GAN, known as EG3D [15], to generate plausible face corrections. The 3D GANs are a type of generative network that is additionally conditioned on camera parameters to generate appearance-consistent facial images under different view angles [15, 16]. To manipulate the camera-to-subject distance, they first estimate an initial head model and camera pose and intrinsics that are jointly optimized through the GAN inversion technique [17] using a combination of the photometric-based loss LPIPS [18] and the L2 loss between the landmark detected on the generated and input face to find the camera parameters, face identity and expressions (i.e. the latent face code) that when provided to the generator produce the distorted input image. The generator is then finetuned (overfit) on this selfie input until convergence. The inversion process is time-consuming as the authors of Disco [7] report that this operation requires more than 2 minutes on a modern GPU. In addition, their pipeline produces subtle-to-noticeable identity changes, depending on the input, due to the inaccuracies of the encoded face from the inversion of 3D GAN outputs. Moreover, similar to [8], their methods require cropping the face from the body leading to additional complex post-processing steps to combine the face with the rest of the body and the background image afterwards.
@@
The work of [20] did not mainly address the task of rectifying perspective distortion, but they were performing it as a subtask of their pipeline which aims to synthesize a selfie photo of the whole body from a group of close-range shots of different body parts. As the face forms a small portion of the final synthesized image, the quality of the corrected face was not the main target of the method. They follow a flow-based warping method similar to LPUP [8] and train the architecture from [21] on image pairs generated by the 3D GAN network EG3D [15]. However, the artefacts in the generated training data led to inferior results compared to previous methods.
@@ III-A Overview
III-B Depth Prediction
We use a residual UNet architecture similar to many recent monocular depth estimation methods [8] for facial depth prediction. The encoder and decoder consist of 5 residual blocks for each. The input to the depth network is the segmented foreground human. Hence, the network only estimates the depth over the foreground region.
III-C Feature Extraction
Instead of directly reprojecting the image in the RGB space, we extract high-level features from the foreground-segmented input image with another UNet-based architecture composed of 5 mirrored pairs of convolution layers. The feature extractor provides an enriched representation of the underlying RGB image to the generation net. It produces a feature map that has the same spatial size as the input but with 64 channels. Thus, every depth pixel is attached to a richer feature vector of length 64.
III-D Image Warping
Once we have the depth map of the subject, we are able to synthesize a novel view of the subject by manipulating the camera-to-subject distance and the focal length. We unproject the face to the 3D space of the original camera $C_{1}$ that has camera intrinsics matrix $K_{1}$ expressed as follows:
@@ III-D Image Warping
Rendering the actual image (i.e. warping the image) is called the rasterization process. To enable E2E training of our model, we leverage a differentiable renderer module of the PyTorch3D framework [5].
@@
III-E Camera Horizontal Translation
@@
To compute the rotation $\theta$, we choose an arbitrary point in the face where this point after translating and rotating the camera by $t_{x}$ and $\theta$ respectively, should have the same position as if we moved the camera back with $t_{x}$ and $\theta$ are zeros. We choose this pixel point on the left face boundary if the subject is closer to the left image side and vice versa. Choosing the point on face boundaries guarantees that this point will not go outside the image boundary after the reprojection. This is especially important when the face is close to the image boundaries; otherwise, any point is valid. Let’s suppose the chosen pixel point $p$ in the input image is at position $(u,v)$ expressed in 2D image coordinates. The 3D coordinates of the point $p$ in camera $C_{2}$ is $P_{2}$ = $X_{2},Y_{2},Z_{2}]^{T}$. The transformation of $P_{2}$ to the coordinates of camera $C^{\prime}_{2}$ is:
@@
We use a ResNet18 [24] architecture for our translation module. It predicts a vector $V$ of length $l=50$ which corresponds to the equal-spacing discretization of the translation range [20,20][-20,20] cm. During inference, we find the translation by computing the Softmax of the output vector as the following:
@@ III-F Image Refinement
III-G Post Processing and Image Composition
@@
III-H Training and Loss Functions
Depth Loss. To train the depth network, we use a loss function that is typically used in training monocular depth estimation networks [8]. Specifically, we use L1 loss (i.e. the mean absolute error) $\mathcal{L}_{L1}$ between the predicted depth and ground truth depth. We also compute $\mathcal{L}_{tanh}$ which is the L1 loss between the hyperbolic tangent (i.e. tanhtanh function) of each of them, and $\mathcal{L}_{grad}$ which is the multi-scale gradients loss [8]. The overall loss for depth training is then defined as the following:
@@ III-H Training and Loss Functions
Horizontal Translation Module Training and Loss. During our experiments, we found it hard to train the translation network (HzT net) in an unsupervised way by incorporating it directly into the pipeline. Thus, we generated the training GT targets to train the network in a supervised way. To compute the GT translations, we discretize the possible range of camera horizontal translation, as mentioned in Section III-E, to 50 translation steps or bins. For each step, we compute the corresponding rotation according to (6), warp the image to the new pose and provide it to our pretrained generator module which predicts the rectified image and mask as explained in Section III-F. We compare the predicted output mask $m_{g}$ with the GT target mask $m_{t}$. The details of synthesizing $m_{t}$ are explained in Section IV. We use the intersection over union (IoU) to compute the difference between $m_{g}$ and $m_{t}$ as the following:
@@
We use the binary cross-entropy loss between the predicted vector $\hat{V}$ from the HzT net and the target vector $V$. The predicted translation is estimated according to (7) and used to warp the input foreground mask $m$. We then compute the IoU loss between the warped mask $m_{w}$ and $m_{t}$. The total loss for training the HzT module becomes as the following:
@@
IV Dataset Generation
To train our model, we need pairs of perspective-distorted images and their distortion-free targets. In [1], they used the BU-4DFE facial dataset [9] for generating training pairs. This dataset has static illumination and expressions for each of the subjects. The 3D head is reconstructed from the synchronized multiple cameras. The reconstruction does not include the subject’s body and usually has limited variations of head poses. The distorted version of the faces are rendered simply using perspective projection and rasterization, without changing illumination or considering complex reflectance properties. Unlike [1], we utilize Unreal Engine (UE) for rendering a realistic synthetic facial dataset. UE offers realistic human characters called Metahuman (MH). Those MH are captured in different locations of a simulated city project in UE called the City Sample. During capturing, we also vary the camera-to-subject distance, the camera orientations, the head pitch, yaw, roll and facial expressions.
@@ IV Dataset Generation
We used two types of MH characters: the first is a predefined set of characters provided by UE, and the second is a subclass instance of the MH character called ‘CitysSampleCrowd’ character. The latter is a pool of different faces, outfits, skin textures, and grooms (i.e. hair, beard, moustache, etc). The character is then randomly generated from those various options. In the first type, we used 66 different characters, from which 10 characters are dedicated for testing. To simulate individuals wearing eyeglasses, we utilize a UE asset package containing 18 different types of sunglasses and reading glasses for both men and women.
We generated two sets of synthetic data: a single-view and a multi-view dataset. In the former, the close-range and its corresponding far-range images are generated. In the multi-view case, we capture multiple far-range views. Specifically, we generate 4 rear views: left, right, centre, and top side of the front camera, as illustrated in Figure 4. The poses of those cameras are fixed relative to the front camera. This dataset is used for the training and evaluation of the horizontal translation module as explained in Section III-H.
@@
To train the Horizontal translation module described in Section III-E, we need to synthesize the ground-truth target image on the fly during training since we have only the view that corresponds to $t_{x}=0$ (i.e. the camera moves straight to the back). To synthesize the target image with $t_{x}\neq 0$, we reproject the view of the rear centre camera to the target pose. Similarly, we warp the other rear views to the same target pose. The missing pixels from the warped centre view are filled from the other reprojected rear views. The synthesized target mask is used during training, while the target colour photo is only used to report the photometric errors.
V Experiments and Results
V-A Training details
@@
Next, we pretrain the feature extractor ($E$) and generator ($G$) nets shown in Figure 1 where the ground truth depth is used in projection instead of the predicted one. The loss function is computed according to (11) where $\mathcal{L}_{depth}$ is disabled, and $\omega_{all}$, $\omega_{face}$, $\omega_{missing}$, $\beta$ and $\rho$ are set to $0.2$, 11, 55, 11 and $20$, respectively. Adam optimizer [29] is used and the learning rate is $5\times 10^{-4}$ and halved every 50 epochs with a total of 200 epochs.
@@ V-B Evaluation metrics
  • Euclidean distance between the estimated facial landmarks of the predicted output and the reference image. We use Mediapipe [31] to estimate dense facial landmarks and use them to align both images by a similarity transform. We then compute the Euclidean distance between the corresponding landmarks.
@@
V-C Results
We compare our pipeline to the following previous portrait correction solutions: Fried et al. [6], LPUP [1], DFN [14], and Disco [7]. Additionally, we compare our method to Shih’s method [4] which is a solution for correcting a different type of face distortion that occurs in images captured with wide-angle cameras when the subject is close to the edges of the image, which is solved by stereographic projection. DualPriors (DP) [32] is a recent method that addresses the same problem as Shih [4].
Since LPUP’s authors [1] do not release their code or datasets, we reimplemented the first part of the LPUP pipeline so that we can compare their proposed pipeline with ours. Specifically, we replicate the training of the Flownet part of the pipeline. The input to the network consists of an RGB image and a depth label. The depth label is repeated to have the same spatial size as the input image and concatenated with the image channels. The network output is a two-channel deformation map used to warp the input image to correct the perspective distortion. Second, we add a differentiable bilinear sampling module to make a backward warping of the GT distortion-free image using the predicted and ground-truth flow maps to the perspective of the input selfie image. The warped images are labelled as fake and real, respectively, and used to train a discriminator network. Thus, the loss function for training the flow map network becomes similar to LPUP [1]. For a fair comparison, we use our synthetically generated dataset for the training. The dataset is preprocessed in the same way as in LPUP. Particularly, we scale and crop the image such that all faces in the images have approximately the same scale, and match the inner corner of each subject’s right eye to a fixed position through all images.
In the qualitative and quantitative comparison with Fried [6] and Disco [7], we are constrained by the results reported in their paper or project websites, as no codes have been released by either method at the time of paper submission.